STMicroelectronics a sensor equipment manufacturer with its factory located in Shenzhen, China. STMicroelectronics products obtained ISO-22301 certification. Assume that STMicroelectronics has recently developed a new model of the sensor after extensive research and development and proved that there is a significant market for the new model. The new model will be put into the market this year and is expected to stay in the market for four years.

Except for the initial investment that will occur immediately, all cash flows will occur at year-end. To make the new model, STMicroelectronics must initially invest US$150 million in production equipment. The equipment can be sold for US$50 million at the end of four years. STMicroelectronics can sell the new model to two distinct markets:

  1. Original manufacturer market—this market consists primarily of the large cell phone companies that buy sensor equipment for their phone. The new model is expected to sell for US$40 per sensor, and the variable cost of each sensor is US$25.

  1. Replacement market—this market consists of all sensor’s equipment purchased after the cell phone has left the factory. The new model in this market is expected to sell for US$50 per sensor, and the variable cost of each sensor is US$25 which is the same as that in the original manufacturer market.

The project will incur US$25 million in marketing and general administrative costs the first year, and this cost is expected to increase at the inflation rate in subsequent years. STMicroelectronics also intends to raise the selling price of the new sensor at the inflation rate. The annual inflation rate is expected to remain constant at 3,50 percent. Variable costs are expected to increase at 1 percent above the inflation rate.

Tech analysts expect cell phone manufacturers will produce 5,5 million new phones this year and production will grow at 2,5 percent per year thereafter. Each new phone needs four sensors. STMicroelectronics expects the new model will capture 10 percent of the original manufacturer market.

Analysts also estimate that the replacement market size will be 15 million sensors this year and that it will grow at 2 percent annually. STMicroelectronics expects the new model will capture 8 percent market share.

The equipment will be depreciated on the straight-line basis. The immediate initial working capital requirement is US$10 million, and the net working capital requirements will be 15 percent of sales. STMicroelectronics's effective corporate tax rate is 40 percent and its required return is 15 percent.

Instruction:

Perform the capital budgeting analysis (NPV, Payback Period, IRR and PI) to determine whether STMicroelectronics should accept or reject the project! Write briefly the necessary assumptions and estimations for your analysis!
Answer #1

Please see the table below step by step. The second column explains how each row has been calculated

Year, N

Linkage

0

1

2

3

4

Original manufacturer market

New phones production (mn)

A

5.50

5.64

5.78

5.92

Annual growth

B

2.50%

2.50%

2.50%

Sensors / phone

C

4.00

4.00

4.00

4.00

Total nos. of sensors

D = A x C

22.00

22.55

23.11

23.69

Market share

E

10.00%

10.00%

10.00%

10.00%

Sensors sold to manufacturer

F = D x E

2.20

2.26

2.31

2.37

Sale price per unit

G

40.00

41.40

42.85

44.35

Annual growth

H = Inflation of 3.5%

3.50%

3.50%

3.50%

Variable cost per unit

I

25.00

26.13

27.30

28.53

Annual growth

J = 1% + inflation 3.5%

4.50%

4.50%

4.50%

Revenue

R1 = F x G

88.00

93.36

99.04

105.07

Variable Cost

VC1 = F x I

55.00

58.91

63.10

67.59

Replacement Market

Market size (mn)

A

15.00

15.30

15.61

15.92

Annual growth

2%

2%

2%

Market share

B

8.00%

8.00%

8.00%

8.00%

Sensors sold in replacement market

C = A x B

1.20

1.22

1.25

1.27

Sale price per unit

D

50.00

51.75

53.56

55.44

Annual growth

H = Inflation of 3.5%

3.50%

3.50%

3.50%

Variable cost per unit

F

25.00

26.13

27.30

28.53

Annual growth

J = 1% + inflation of 3.5%

4.50%

4.50%

4.50%

Revenue

R2 = C x D

60.00

63.34

66.87

70.59

Variable Cost

VC2 = C x F

30.00

31.98

34.08

36.33

Total Revenue

R = R1 + R2

148.00

156.70

165.91

175.66

[-] Total variable Cost

VC = VC1 + VC2

85.00

90.89

97.19

103.92

Gross Margin

GM = R - VC

63.00

65.81

68.72

71.74

[-] S, G & A Cost

FC

25.00

25.88

26.78

27.72

Annual escalation

= inflation rate of 3.5%

3.50%

3.50%

3.50%

[-] Depreciation

25% x I

37.50

37.50

37.50

37.50

Operating Income

EBIT

0.50

2.44

4.44

6.53

NOPAT

K =EBIT x (1-tax rate of 40%)

0.30

1.46

2.67

3.92

[+] Depreciation

37.50

37.50

37.50

37.50

Operating cash flow

OCF

37.80

38.96

40.17

41.42

Initial investment

I

(150.00)

Working capital requirement

15% x R

10.00

23.50

24.89

26.35

-

Working capital investment

W

(10.00)

(13.50)

(1.38)

(1.46)

26.35

Salvage value at the end of 4 years

S

50

Taxable base at the end of 4 years

B

-

Profit on sale

P = S - B

50.00

[-] Taxes on profit

T = tax rate of 40% x P

20.00

Post tax salvage value

PS = S - T

30.00

Net cash flows related to investment

N = I + W + PS

(160.00)

(13.50)

(1.38)

(1.46)

56.35

Net cash flows

OCF + N

(160.00)

24.30

37.58

38.70

97.76

Cumulative net cash flows (160.00) (135.70) (98.13) (59.42) 38.34

PV factor at 15% discount rate

L

1.0000

0.8696

0.7561

0.6575

0.5718

PV of Net cash flows

M = N x L

(160.00)

21.13

28.42

25.45

55.90

NPV

(29.11)

IRR

Using excel function

7%

PI

(NPV + Initial investment)/ Initial investment

0.82

NPV is negative

IRR = 7% less than required rate of return of 15%

PI = 0.82 i.e. less than 1

Based on all these capital budgeting tools, the project should be rejected.

Payback period: Please see the row containing cumulative cash flows that become positive for the first time after year 3 but before year 4

Hence, payback period will be an interpolated value between year 3 & 4 to be approximated by = 3 + 59.42 / 97.76 = 3.61 years

Necessary assumptions:

1. Initial investment related cash flows occur at the beginning of the period. Rest all cash flows occur at the end of the period.

2. Net working capital in period 1 = 15% of sales in period 2 and so on.

3. Payback period has been estimated by interpolation between year 3 and 4 assuming cashflow during year 4 is uniformly distributed during the year.